OpenStudy (anonymous):

lim x->0, (sin 9x)/(x cos x),,, can someone help?

OpenStudy (zarkon):

9

OpenStudy (zarkon):

write \[\frac{\sin(9x)}{x}\] as \[9\frac{\sin(9x)}{9x}\] \[\frac{\sin(9x)}{9x}\to 1\text{ as }x\to 0\]

OpenStudy (anonymous):

where does the cos x go?

OpenStudy (zarkon):

\[\cos(x)\to 1\text{ as }x\to 0\]

OpenStudy (zarkon):

\[\cos(0)=1\]

OpenStudy (anonymous):

i didnt get it,, u substitute x with 0 is it?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{\sin(9x)}{x}\times \lim_{x\rightarrow 0}\frac{1}{\cos(x)}\] \[\lim_{x \rightarrow 0}\frac{9\sin(9x)}{9x}\times \lim_{x\rightarrow 0}\frac{1}{\cos(x)}\] \[9\lim_{x \rightarrow 0}\frac{\sin(9x)}{9x}\times \lim_{x\rightarrow 0}\frac{1}{\cos(x)}\] first and second limit is 1 the first because you know \[\lim_{x \rightarrow 0}\frac{\sin(x)}{x}=1\] the second by replacing x by 1 (since it is continuous no problem doing that)

OpenStudy (anonymous):

yes! tanx :D

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

MEDAL PLSS

OpenStudy (anonymous):

medal for wht? u dont even answer my ques :D

OpenStudy (agreene):

You could also use L'Hopitals. \[\lim_{x\rightarrow0} \frac{9\cos(9x)}{\cos(x)-x\sin(x)}\]